# A SAT-centered XAI method for Deep Learning based Video Understanding

### v1

@@ 3.2 Formal Representation of Deep Learning Models for Video
The first crucial step in the proposed method is to develop a formal way to represent the deep learning model used for video understanding within a logical framework [62]. This involves encoding both the architecture and the parameters of the neural network using propositional logic [65].
@@ 3.2.1 Encoding Neural Network Architectures and Parameters in a Logical Framework
One possible approach to encoding a deep learning model in propositional logic involves discretizing the continuous activations and weights of the network [48, 72]. For a given pre-trained deep learning model, each neuron’s activation and each weight connecting neurons can be represented by a set of Boolean variables. The number of Boolean variables used to represent each continuous value would determine the precision of the discretization.
For example, a real-valued activation $a$ could be represented by $n$ Boolean variables $b_{1},b_{2},\dots,b_{n}$, where the binary sequence $b_{n}b_{n-1}\dots b_{1}$ represents a quantized version of $a$ within a certain range and precision. Similarly, weights and biases can also be discretized and represented using Boolean variables [73].
@@
where $f$ is the activation function, $w_{ij}^{(l)}$ is the weight connecting neuron $i$ in layer $l-1$ to neuron $j$ in layer $l$, and $b_{j}^{(l)}$ is the bias [39, 46].
@@
Consider a simplified example: a neuron with two binary inputs $x_{1},x_{2}$, weights $w_{1}=1,w_{2}=1$, bias $b=0$, and a threshold activation function that outputs 1 if the weighted sum is $\geq 2$ and 0 otherwise. The weighted sum is $x_{1}+x_{2}$. The output is 1 only if $x_{1}=1$ and $x_{2}=1$. This can be directly represented by the clause $(x_{1}\wedge x_{2})\Leftrightarrow y$, which in CNF is ($(\neg x_{1}\vee\neg x_{2}\vee y)\wedge(x_{1}\vee\neg y)\wedge(x_{2}\vee\neg y)$ [40, 45].
For more complex scenarios with discretized real values, the encoding would involve a larger number of Boolean variables and clauses to represent the arithmetic operations and the activation functions with sufficient precision. The level of abstraction for the encoding can also be adjusted. For instance, instead of encoding individual neurons, one could potentially encode the overall input-output behavior of certain layers or modules using logical rules learned from the trained network [44].
@@ 3.2.2 Handling Spatio-Temporal Dependencies in Video Data
Video data introduces the challenge of representing both spatial information within each frame and the temporal relationships between consecutive frames. In the proposed method, the input video would first be pre-processed to extract relevant features. These features could be high-level semantic features obtained from pre-trained models or lower-level features like motion vectors or visual descriptors [76].
@@
The temporal connections between frames could be handled by explicitly representing the flow of information through the deep learning model over time. For models like RNNs or LSTMs, the state transitions between time steps would need to be encoded as logical clauses. For models like 3D CNNs or Transformer networks that process the entire video or chunks of it at once, the spatial and temporal relationships are implicitly learned within the network’s architecture. The encoding would then need to capture these learned relationships in the logical framework [17, 64].
@@ 3.3 SAT-Based Query Formulation for Explainability
Once the deep learning model and the video data are represented within the logical framework, the next step is to formulate the explanation queries as satisfiability problems. This involves adding logical constraints to the encoded system that correspond to the specific question being asked [4].
@@ 3.3.1 Defining Explanation Queries as Satisfiability Problems
Both ”Why?” (abductive) and ”Why not?” (contrastive) questions can be formulated as SAT problems within this framework [7, 11].
@@ 3.3.2 Formulating ”Why?” and ”Why Not?” Questions using Logical Constraints
The explanation queries are formalized by adding specific logical constraints to the SAT problem. These constraints depend on the type of explanation being sought [31, 43].
@@
We then aim to find a minimal $F$ for which this problem is satisfiable [3, 5].
@@
where $V^{\prime}$ represents a modified version of the input video. The goal is to find a $V^{\prime}$ that satisfies this formula and minimizes the difference between $V$ and $V^{\prime}$ according to some distance metric. This difference would represent the minimal change needed to flip the prediction to $c_{2}$ [9, 56]. The distance metric can be incorporated into the SAT problem by introducing additional variables and clauses that penalize changes to the original input features.
@@ 3.4 Novel Model Architecture: Integrating a SAT Solver with a Deep Learning Video Understanding Model
The proposed model architecture integrates a pre-trained deep learning model for video understanding with a SAT solver to generate formal explanations [38, 19].
@@ 3.4.1 Detailed Description of the Architecture and its Components
    Pre-trained Deep Learning Model for Video Understanding: This is the model whose predictions we want to explain. It could be any state-of-the-art model suitable for the video understanding task at hand (e.g., a 3D CNN for action recognition) [63, 30].
@@
    Encoding Module: This module takes the pre-trained deep learning model and an input video as input and translates them into a SAT problem represented as a CNF formula. This involves discretizing the model’s weights, biases, and activations, as well as the features extracted from the video, and representing the model’s operations and the video data using Boolean variables and logical clauses [61, 21].
@@
    Query Formulation Module: This module takes an explanation query (e.g., ”Why this prediction?”, ”Why not that prediction?”) and formulates it as a set of logical constraints that are added to the SAT problem generated by the encoding module [28, 29].
@@
    SAT Solver: A high-performance SAT solver (e.g., a CDCL solver) is used to find a satisfying assignment to the complete SAT problem (encoded model + encoded video + query constraints) [35, 32].
@@
    Explanation Extraction Module: This module takes the output of the SAT solver (a satisfying assignment or an indication of unsatisfiability) and interprets it back in the context of the original deep learning model and video data to generate a formal explanation. The explanation could be in the form of a minimal set of necessary input features, a minimal change to the input that would flip the prediction, or a set of logical rules that justify the prediction [36, 24].
The flow of information in this architecture is as follows: First, an input video is fed into the encoding module along with the pre-trained deep learning model. The encoding module generates a CNF formula representing the model and the video. Then, an explanation query is provided to the query formulation module, which adds logical constraints to the CNF formula based on the query. This complete SAT problem is then passed to the SAT solver. The SAT solver attempts to find a satisfying assignment. If a satisfying assignment is found, the explanation extraction module interprets this assignment to produce the explanation. If the SAT solver determines that the problem is unsatisfiable, it might indicate that no explanation of the requested type exists within the defined framework [20, 37].
@@ 3.4.2 Workflow of the Explanation Generation Process
    Input: A pre-trained deep learning model for video understanding, an input video, and an explanation query [2, 69].
@@
    Encoding: The encoding module discretizes the weights, biases, and activations of the deep learning model. It also extracts features from the input video and discretizes them. Both the model’s operations and the video features are then translated into a CNF formula using Boolean variables and logical clauses [34].
@@ 4.1 Technical and Mathematical Details
This section delves into the technical and mathematical aspects of the proposed method, providing formal definitions, details of the logical encodings, and algorithms for explanation generation [78].
@@ 4.1.1 Formal Definitions of Key Concepts
  • Discretized Neural Network State: The state of a deep learning model (weights, biases, activations) is represented by a set of Boolean variables obtained through a discretization process. Let $W$ be the set of weights, $B$ the set of biases, and $A$ the set of activations in the original model. Their discretized counterparts are represented by sets of Boolean variables $B_{W}$, $B_{B}$, $B_{A}$ [79].
  • Discretized Video Features: Features extracted from the input video at different time steps are also discretized and represented by a set of Boolean variables $B_{F}$ [70].
  • Logical Encoding Function: A function $E$ that maps the deep learning model and the input video to a CNF formula $\phi=E(\text{Model},\text{Video})$ over the Boolean variables $B_{W}\cup B_{B}\cup B_{A}\cup B_{F}\cup B_{O}$, where $B_{O}$ represents the discretized output of the model [47].
  • Explanation Query Constraints: A set of logical clauses $Q$ that represent the specific explanation query (e.g., target prediction, constraints on input features) [13].
@@ 4.1.2 Logical Encodings and Transformations
The logical encoding function $E$ needs to capture the behavior of each layer and operation in the deep learning model [71, 14]. For a linear layer with weights $w_{ij}$ and inputs $x_{i}$, the output $y_{j}=\sum_{i}w_{ij}x_{i}$ is first computed using discretized representations of $w_{ij}$ and $x_{i}$. The multiplication and addition of discretized values can be implemented using Boolean logic.
For example, consider the multiplication of two discretized values. If $x$ and $w$ are each represented by $n$ bits, the multiplication can be implemented using a series of bitwise AND and shift operations, similar to how multiplication is done in digital circuits. Each bit of the product can be represented by a Boolean variable, and the relationships between these variables and the input bits can be expressed as CNF clauses. Addition can be implemented using full adders, where each bit of the sum and the carry-out bit are represented by Boolean variables, and their relationships with the input bits and carry-in bit are encoded as CNF clauses [15, 41].
The activation function applied to the output of the linear layer also needs to be encoded. For a ReLU function ($\max(0,z)$), if $z$ is represented by $n$ Boolean variables, the output will be 0 if $z\leq 0$ and $z$ if $z>0$. This can be encoded by checking the sign bit of the discretized $z$. If the sign bit indicates negative, the output Boolean variables are all set to false (representing 0). Otherwise, they are set to the same values as the Boolean variables representing $z$ [27, 62]. Similar encodings can be developed for other activation functions using piecewise linear approximations or by directly implementing their logical equivalents on the discretized representations.
For video data, if features are extracted as real-valued vectors at each time step, each element of these vectors is discretized into a set of Boolean variables. The temporal relationships, if modeled explicitly (e.g., in an RNN), would require encoding the state transitions between time steps. This would involve representing the hidden state at each time step using Boolean variables and encoding the update rules of the RNN (e.g., for LSTM, the update of the cell state and hidden state) using logical clauses based on the discretized inputs, previous hidden state, and weights [65, 48].
@@ 4.1.4 Mathematical Proofs
Proving the soundness and completeness of this approach depends heavily on the accuracy of the logical encoding of the deep learning model and the video data [9, 56].
@@
If Algorithm 1 finds a set of necessary input features $E$, and the SAT solver returns a satisfying assignment when only these features are fixed (along with the model and the target prediction), then it implies that the model’s internal logic, as captured by the encoding, can still lead to the target prediction based solely on these features. The encoding aims to be a faithful representation of the discretized model’s behavior [38, 19].
@@
Completeness would imply that if there exists a minimal set of necessary features (or minimal changes for a contrastive explanation), Algorithm 1 (or 2) will find it. This depends on the SAT solver’s ability to find a satisfying assignment if one exists and on the effectiveness of the iterative minimization process in Algorithm 1 or the optimization process in Algorithm 2. For Algorithm 1, the iterative removal of features ensures that the resulting set is minimal with respect to the features considered. For Algorithm 2, if the SAT solver with the optimization objective finds a solution, it will be one with the minimum number of changes within the limitations of the encoding and the solver’s capabilities [63, 30].
Formal proofs would require a more rigorous definition of the encoding function $E$ and the relationship between the continuous model and its discrete logical representation. This could involve bounding the error introduced by discretization and showing that the logical clauses correctly implement the intended arithmetic and logical operations [61, 21].
@@ 5.1 In-depth Discussion of the Proposed Method and its Implications
The proposed method offers a novel approach to explainability in deep learning for video understanding by leveraging the rigor of formal methods and the power of SAT solving [78].
@@ 5.1.1 Advantages over Existing XAI Techniques for Video Understanding
Compared to many existing XAI techniques for video understanding, such as spatio-temporal attention or gradient-based saliency maps, the proposed method has the potential to provide explanations with formal guarantees [79]. While visual explanations can be intuitive, they often lack a clear logical justification for the model’s prediction. The SAT-based approach, by encoding the model’s behavior and the explanation query in a logical framework, can provide explanations that are rooted in logical deduction [70, 47]. For ”Why?” explanations, the identified necessary input features are those that, according to the encoded model’s logic, are essential for achieving the prediction [13]. For ”Why not?” explanations, the minimal changes found are those that, when applied to the input, satisfy the logical constraints of the model predicting the alternative outcome. This contrasts with heuristic methods that might highlight features that are merely correlated with the prediction rather than being logically necessary or sufficient [71, 14].
Furthermore, the proposed method can potentially address the limitations of current techniques in explaining long-term temporal reasoning in videos [15, 41]. By encoding the temporal dynamics of the deep learning model (e.g., through the encoding of recurrent layers or by considering the entire spatio-temporal input for models like 3D CNNs), the SAT solver can reason about how events across different time steps contribute to the final prediction [27, 62]. This could lead to more comprehensive explanations for video-level tasks that involve understanding sequences of actions or events over time [65].

### v2

@@ 3.2 Formal Representation of Deep Learning Models for Video
The first crucial step in the proposed method is to develop a formal way to represent the deep learning model used for video understanding within a logical framework. This involves encoding both the architecture and the parameters of the neural network using propositional logic.
@@ 3.2.1 Encoding Neural Network Architectures and Parameters in a Logical Framework
One possible approach to encoding a deep learning model in propositional logic involves discretizing the continuous activations and weights of the network. For a given pre-trained deep learning model, each neuron’s activation and each weight connecting neurons can be represented by a set of Boolean variables. The number of Boolean variables used to represent each continuous value would determine the precision of the discretization.
For example, a real-valued activation $a$ could be represented by $n$ Boolean variables $b_{1},b_{2},\dots,b_{n}$, where the binary sequence $b_{n}b_{n-1}\dots b_{1}$ represents a quantized version of $a$ within a certain range and precision. Similarly, weights and biases can also be discretized and represented using Boolean variables.
@@
where $f$ is the activation function, $w_{ij}^{(l)}$ is the weight connecting neuron $i$ in layer $l-1$ to neuron $j$ in layer $l$, and $b_{j}^{(l)}$ is the bias.
@@
Consider a simplified example: a neuron with two binary inputs $x_{1},x_{2}$, weights $w_{1}=1,w_{2}=1$, bias $b=0$, and a threshold activation function that outputs 1 if the weighted sum is $\geq 2$ and 0 otherwise. The weighted sum is $x_{1}+x_{2}$. The output is 1 only if $x_{1}=1$ and $x_{2}=1$. This can be directly represented by the clause $(x_{1}\wedge x_{2})\Leftrightarrow y$, which in CNF is ($(\neg x_{1}\vee\neg x_{2}\vee y)\wedge(x_{1}\vee\neg y)\wedge(x_{2}\vee\neg y)$.
For more complex scenarios with discretized real values, the encoding would involve a larger number of Boolean variables and clauses to represent the arithmetic operations and the activation functions with sufficient precision. The level of abstraction for the encoding can also be adjusted. For instance, instead of encoding individual neurons, one could potentially encode the overall input-output behavior of certain layers or modules using logical rules learned from the trained network.
@@ 3.2.2 Handling Spatio-Temporal Dependencies in Video Data
Video data introduces the challenge of representing both spatial information within each frame and the temporal relationships between consecutive frames. In the proposed method, the input video would first be pre-processed to extract relevant features. These features could be high-level semantic features obtained from pre-trained models or lower-level features like motion vectors or visual descriptors.
@@
The temporal connections between frames could be handled by explicitly representing the flow of information through the deep learning model over time. For models like RNNs or LSTMs, the state transitions between time steps would need to be encoded as logical clauses. For models like 3D CNNs or Transformer networks that process the entire video or chunks of it at once, the spatial and temporal relationships are implicitly learned within the network’s architecture. The encoding would then need to capture these learned relationships in the logical framework.
@@ 3.3 SAT-Based Query Formulation for Explainability
Once the deep learning model and the video data are represented within the logical framework, the next step is to formulate the explanation queries as satisfiability problems. This involves adding logical constraints to the encoded system that correspond to the specific question being asked.
@@ 3.3.1 Defining Explanation Queries as Satisfiability Problems
Both ”Why?” (abductive) and ”Why not?” (contrastive) questions can be formulated as SAT problems within this framework.
@@ 3.3.2 Formulating ”Why?” and ”Why Not?” Questions using Logical Constraints
The explanation queries are formalized by adding specific logical constraints to the SAT problem. These constraints depend on the type of explanation being sought.
@@
We then aim to find a minimal $F$ for which this problem is satisfiable.
@@
where $V^{\prime}$ represents a modified version of the input video. The goal is to find a $V^{\prime}$ that satisfies this formula and minimizes the difference between $V$ and $V^{\prime}$ according to some distance metric. This difference would represent the minimal change needed to flip the prediction to $c_{2}$. The distance metric can be incorporated into the SAT problem by introducing additional variables and clauses that penalize changes to the original input features.
@@ 3.4 Novel Model Architecture: Integrating a SAT Solver with a Deep Learning Video Understanding Model
The proposed model architecture integrates a pre-trained deep learning model for video understanding with a SAT solver to generate formal explanations.
@@ 3.4.1 Detailed Description of the Architecture and its Components
    Pre-trained Deep Learning Model for Video Understanding: This is the model whose predictions we want to explain. It could be any state-of-the-art model suitable for the video understanding task at hand (e.g., a 3D CNN for action recognition).
@@
    Encoding Module: This module takes the pre-trained deep learning model and an input video as input and translates them into a SAT problem represented as a CNF formula. This involves discretizing the model’s weights, biases, and activations, as well as the features extracted from the video, and representing the model’s operations and the video data using Boolean variables and logical clauses.
@@
    Query Formulation Module: This module takes an explanation query (e.g., ”Why this prediction?”, ”Why not that prediction?”) and formulates it as a set of logical constraints that are added to the SAT problem generated by the encoding module.
@@
    SAT Solver: A high-performance SAT solver (e.g., a CDCL solver) is used to find a satisfying assignment to the complete SAT problem (encoded model + encoded video + query constraints).
@@
    Explanation Extraction Module: This module takes the output of the SAT solver (a satisfying assignment or an indication of unsatisfiability) and interprets it back in the context of the original deep learning model and video data to generate a formal explanation. The explanation could be in the form of a minimal set of necessary input features, a minimal change to the input that would flip the prediction, or a set of logical rules that justify the prediction.
The flow of information in this architecture is as follows: First, an input video is fed into the encoding module along with the pre-trained deep learning model. The encoding module generates a CNF formula representing the model and the video. Then, an explanation query is provided to the query formulation module, which adds logical constraints to the CNF formula based on the query. This complete SAT problem is then passed to the SAT solver. The SAT solver attempts to find a satisfying assignment. If a satisfying assignment is found, the explanation extraction module interprets this assignment to produce the explanation. If the SAT solver determines that the problem is unsatisfiable, it might indicate that no explanation of the requested type exists within the defined framework.
@@ 3.4.2 Workflow of the Explanation Generation Process
    Input: A pre-trained deep learning model for video understanding, an input video, and an explanation query.
@@
    Encoding: The encoding module discretizes the weights, biases, and activations of the deep learning model. It also extracts features from the input video and discretizes them. Both the model’s operations and the video features are then translated into a CNF formula using Boolean variables and logical clauses.
@@ 4.1.1 Formal Definitions of Key Concepts
  • Discretized Neural Network State: The state of a deep learning model (weights, biases, activations) is represented by a set of Boolean variables obtained through a discretization process. Let $W$ be the set of weights, $B$ the set of biases, and $A$ the set of activations in the original model. Their discretized counterparts are represented by sets of Boolean variables $B_{W}$, $B_{B}$, $B_{A}$.
  • Discretized Video Features: Features extracted from the input video at different time steps are also discretized and represented by a set of Boolean variables $B_{F}$.
  • Logical Encoding Function: A function $E$ that maps the deep learning model and the input video to a CNF formula $\phi=E(\text{Model},\text{Video})$ over the Boolean variables $B_{W}\cup B_{B}\cup B_{A}\cup B_{F}\cup B_{O}$, where $B_{O}$ represents the discretized output of the model.
  • Explanation Query Constraints: A set of logical clauses $Q$ that represent the specific explanation query (e.g., target prediction, constraints on input features).
@@ 4.1.2 Logical Encodings and Transformations
The logical encoding function $E$ needs to capture the behavior of each layer and operation in the deep learning model. For a linear layer with weights $w_{ij}$ and inputs $x_{i}$, the output $y_{j}=\sum_{i}w_{ij}x_{i}$ is first computed using discretized representations of $w_{ij}$ and $x_{i}$. The multiplication and addition of discretized values can be implemented using Boolean logic.
For example, consider the multiplication of two discretized values. If $x$ and $w$ are each represented by $n$ bits, the multiplication can be implemented using a series of bitwise AND and shift operations, similar to how multiplication is done in digital circuits. Each bit of the product can be represented by a Boolean variable, and the relationships between these variables and the input bits can be expressed as CNF clauses. Addition can be implemented using full adders, where each bit of the sum and the carry-out bit are represented by Boolean variables, and their relationships with the input bits and carry-in bit are encoded as CNF clauses.
The activation function applied to the output of the linear layer also needs to be encoded. For a ReLU function ($\max(0,z)$), if $z$ is represented by $n$ Boolean variables, the output will be 0 if $z\leq 0$ and $z$ if $z>0$. This can be encoded by checking the sign bit of the discretized $z$. If the sign bit indicates negative, the output Boolean variables are all set to false (representing 0). Otherwise, they are set to the same values as the Boolean variables representing $z$. Similar encodings can be developed for other activation functions using piecewise linear approximations or by directly implementing their logical equivalents on the discretized representations.
For video data, if features are extracted as real-valued vectors at each time step, each element of these vectors is discretized into a set of Boolean variables. The temporal relationships, if modeled explicitly (e.g., in an RNN), would require encoding the state transitions between time steps. This would involve representing the hidden state at each time step using Boolean variables and encoding the update rules of the RNN (e.g., for LSTM, the update of the cell state and hidden state) using logical clauses based on the discretized inputs, previous hidden state, and weights.
@@ 4.1.4 Mathematical Proofs
Proving the soundness and completeness of this approach depends heavily on the accuracy of the logical encoding of the deep learning model and the video data.
@@
If Algorithm 1 finds a set of necessary input features $E$, and the SAT solver returns a satisfying assignment when only these features are fixed (along with the model and the target prediction), then it implies that the model’s internal logic, as captured by the encoding, can still lead to the target prediction based solely on these features. The encoding aims to be a faithful representation of the discretized model’s behavior.
@@
Completeness would imply that if there exists a minimal set of necessary features (or minimal changes for a contrastive explanation), Algorithm 1 (or 2) will find it. This depends on the SAT solver’s ability to find a satisfying assignment if one exists and on the effectiveness of the iterative minimization process in Algorithm 1 or the optimization process in Algorithm 2. For Algorithm 1, the iterative removal of features ensures that the resulting set is minimal with respect to the features considered. For Algorithm 2, if the SAT solver with the optimization objective finds a solution, it will be one with the minimum number of changes within the limitations of the encoding and the solver’s capabilities.
Formal proofs would require a more rigorous definition of the encoding function $E$ and the relationship between the continuous model and its discrete logical representation. This could involve bounding the error introduced by discretization and showing that the logical clauses correctly implement the intended arithmetic and logical operations.
@@ 5.1 In-depth Discussion of the Proposed Method and its Implications
The proposed method offers a novel approach to explainability in deep learning for video understanding by leveraging the rigor of formal methods and the power of SAT solving.
@@ 5.1.1 Advantages over Existing XAI Techniques for Video Understanding
Compared to many existing XAI techniques for video understanding, such as spatio-temporal attention or gradient-based saliency maps, the proposed method has the potential to provide explanations with formal guarantees. While visual explanations can be intuitive, they often lack a clear logical justification for the model’s prediction. The SAT-based approach, by encoding the model’s behavior and the explanation query in a logical framework, can provide explanations that are rooted in logical deduction. For ”Why?” explanations, the identified necessary input features are those that, according to the encoded model’s logic, are essential for achieving the prediction. For ”Why not?” explanations, the minimal changes found are those that, when applied to the input, satisfy the logical constraints of the model predicting the alternative outcome. This contrasts with heuristic methods that might highlight features that are merely correlated with the prediction rather than being logically necessary or sufficient.
Furthermore, the proposed method can potentially address the limitations of current techniques in explaining long-term temporal reasoning in videos. By encoding the temporal dynamics of the deep learning model (e.g., through the encoding of recurrent layers or by considering the entire spatio-temporal input for models like 3D CNNs), the SAT solver can reason about how events across different time steps contribute to the final prediction. This could lead to more comprehensive explanations for video-level tasks that involve understanding sequences of actions or events over time.